OpenStudy (anonymous):

So I know that the lim (f(x) + g(x)) = lim f(x) + lim g(x). I was wondering if the converse is true. If so, how do I prove it. Thanks!

OpenStudy (anonymous):

Yes, there is a way to prove that. Can I show you by an example?

OpenStudy (anonymous):

Sure. Thanks.

OpenStudy (anonymous):

now let's do the left side

OpenStudy (anonymous):

Thanks. That makes sense. Do you know if there is a formal proof for the converse? I know the proof for the one above, I'm just stumped on how to set up the proof for the converse.

OpenStudy (anonymous):

Kind of. I was wondering if there was a formal proof for this?

OpenStudy (anonymous):

Like is there a delta/episilon proof for the converse or some other type of proof for the converse. I know the delta/episilon proof for the original, just not sure how to approach the converse.

OpenStudy (anonymous):

I'll keep looking out for more formal proofs of this limit. I'll message you if I find more.

OpenStudy (anonymous):

I've looked all over the internet and can't find anything. Thanks for your help!

OpenStudy (anonymous):

Daw, but there are some peeps on Openstudy that is very resourceful. You can try asking them too.

OpenStudy (anonymous):

or him*

OpenStudy (anonymous):

Thanks, I appreciate your help.

OpenStudy (anonymous):

He might be busy, sleeping I think. I can try asking him later. He's the guy that always have a lot of good resources.

OpenStudy (anonymous):

You're welcomed :) see you around soon.

OpenStudy (anonymous):

thank you @xapproachesinfinity

OpenStudy (anonymous):

@xapproachesinfinity the top is me proving it wrong. it's actually involved episilon and delta to prove the limit, which is beyond my level of understanding. from what I know this the college level of calc, and he's a math major taking his college final exam in the next friday.

OpenStudy (anonymous):

Btw, your profile pic looks really yummy

OpenStudy (xapproachesinfinity):

hmm, I'm not sure how to go around the delta epsilon proof. the delta epsilon states that if |x-a|<delta then |f(x)-L|<epsilon may be you need to set up this definition to two functions to prove the converse.

OpenStudy (xapproachesinfinity):

haha that is called tagine lol

OpenStudy (xapproachesinfinity):

what are your thoughts? eru?

OpenStudy (anonymous):

Thanks xapproach :) Idk him well but I thought this is important because it would affect his math major career so. Thank you so much for coming in. I tried asking alot of people >.<

OpenStudy (xapproachesinfinity):

I guess i haven't seen a poof to this before! I'm kinda rusty with proofs now lol. can we set up something to work with like let L1=lim f as x goes to c and L2=lim g as x goes to c

OpenStudy (anonymous):

hm, do you think this is it? I don't know how to read this type of proof so I'm not sure if this is converse or not. http://math.stackexchange.com/questions/408135/prove-limsup-limits-n-to-infty-a-nb-n-le-limsup-limits-n-to-infty?rq=1

OpenStudy (anonymous):

especially when they are using a sub n , b sub n not f(x) g(x)

OpenStudy (xapproachesinfinity):

those are supremum proof! I haven't studied that as far as i remember. that's dealing with ap, gp

OpenStudy (xapproachesinfinity):

may be the property |a+b|<|a|+|b| will be of a help here! not sure

OpenStudy (anonymous):

oh I have ZERO clue what episilon mean xD left alone ap , gp. You're already really good! This type of math is just over our knowledge >.<

OpenStudy (xapproachesinfinity):

Arithmetic progression and geometric progression. you're not familiar with those

OpenStudy (xapproachesinfinity):

@dan815 help here please!

OpenStudy (anonymous):

oh so that's what it means. Yes I'm familiar, but the calc I'm dealing with doesn't include using episilon so. so we just give up? : o

OpenStudy (xapproachesinfinity):

hehe, there sure will be someone to help! I familiar with thee epsilon definition but I forget what i knew lol. that's why i'm starting over

OpenStudy (anonymous):

@SithsAndGiggles please help with this!

OpenStudy (anonymous):

oh, good luck! I hope someone finds the answer to this v.v tried for 4 hours already xD

OpenStudy (anonymous):

thanks again xapproaches. You're so nice :))

OpenStudy (xapproachesinfinity):

I haven't done anything lol, but anytime!

OpenStudy (xapproachesinfinity):

I want to see how to get the answer to this too now haha

OpenStudy (xapproachesinfinity):

|x-c|<deltha1==>|f(x)-L1|<epsilon1 |x-c|<deltha2==>|f(x)-L2|<epsilon2 can we use any property here?

OpenStudy (xapproachesinfinity):

We assume that L1 and L2 exist for f and g respectively

OpenStudy (xapproachesinfinity):

the second line I meant g(x)

OpenStudy (anonymous):

If we try, we'll get it eventually : D

OpenStudy (anonymous):

so far I'm understanding what you're proving :D

OpenStudy (xapproachesinfinity):

This from some of limits to limit of sum so it the converse? that's what we are looking for

OpenStudy (xapproachesinfinity):

Oh good! I need a hand for this lol

OpenStudy (xapproachesinfinity):

can we add the two statements? not sure if it is correct though

OpenStudy (xapproachesinfinity):

like 2|x-c|<deltha1 +deltha 2 ==> |f(x)-L1|+|g(x)-L2|< epsilon1 + epislon2 this weird not sure if it is logical

OpenStudy (xapproachesinfinity):

may be 2|x-c|<deltha1 + deltha2 is good to verify but the implication we need someone to check this haha

OpenStudy (anonymous):

oh god xD sometimes I really hate myself for not knowing things. Let me tag one more person to check--- yes, we should ask!

OpenStudy (anonymous):

maybe nincompoop?

OpenStudy (xapproachesinfinity):

Tagg him/her!

OpenStudy (anonymous):

@nincompoop please help check this proof? Thank you Nin

OpenStudy (anonymous):

Really hope he comes.

OpenStudy (xapproachesinfinity):

|f(x)-L1|+|g(x)-L2|> |f(x)+g(x)-(L1+L2)|

OpenStudy (xapproachesinfinity):

this seem the flaw little bit haha

OpenStudy (xapproachesinfinity):

Now we got |f(x)+g(x)-(L1+L2)|<epsilon1 + epsilon2

OpenStudy (xapproachesinfinity):

we are half way now?

OpenStudy (xapproachesinfinity):

@ganeshie8 can you check this please! I'm sure I'm doin' this right

OpenStudy (xapproachesinfinity):

not*

OpenStudy (anonymous):

I think we are xD but not sure if we should go further

OpenStudy (anonymous):

until someone comes and checks

OpenStudy (anonymous):

@tkhunny please help check this proof, we are so unsure of its accuracy

OpenStudy (xapproachesinfinity):

if |x-c|<(deltha1+deltha2)/2 holds as we said above then it implies |(f(x)+g(x))-(L1+L2)|<epsilon1+epsilon2 can we deduce that the limit of (f(x)+g(x)) exist and it is L1+L2? or may be we need further steps! based on the epsilon definition this should be good, assuming the above step are true!

OpenStudy (anonymous):

Yes! All we need is someone clarifying this.

OpenStudy (anonymous):

Thanks for your help!

OpenStudy (anonymous):

Just curious, ehub, how did you do it?

OpenStudy (anonymous):

and is our way of doing this proof correct?

OpenStudy (xapproachesinfinity):

@ehubbard0307, you are welcome! not sure this is the answer? please wait for ppl who have more resources on this

OpenStudy (xapproachesinfinity):

looks fine to me?

OpenStudy (xapproachesinfinity):

@ash2326

OpenStudy (anonymous):

@ehubbard0307 have you got your answer to this question yet? If yes, I hope you don't mind sharing it with us, or perhaps check our proof? I'm not sure of what your message means >.<

OpenStudy (xapproachesinfinity):

@aum, see this please!

OpenStudy (xapproachesinfinity):

I'm coming back!

OpenStudy (anonymous):

Ok!

OpenStudy (xapproachesinfinity):

the sleep will not come if this is left unsolved haha

OpenStudy (tkhunny):

Consider f(x) = x and g(x) = -x

OpenStudy (anonymous):

Assuming\[\begin{cases}\lim_{x\to c}f(x)=L\\\lim_{x\to c}g(x)=M\end{cases}\] you want to show that \[\lim_{x\to c}\bigg(f(x)+g(x)\bigg)=L+M\] For the proof, you want to find \(\delta\) such that \(0<|x-c|<\delta\) implies \(\left|\bigg(f(x)+g(x)\bigg)-\bigg(L+M\bigg)\right|<\epsilon\). To do that, you work backwards with the previous inequality and rewrite as needed: \[\begin{align*}\left|\bigg(f(x)+g(x)\bigg)-\bigg(L+M\bigg)\right|&=\left|\bigg(f(x)-L\bigg)+\bigg(g(x)-M\bigg)\right|&(1)\\ &\le |f(x)-L|+|g(x)-M|&(2)\\ &<\frac{\epsilon}{2}+\frac{\epsilon}{2}&(3)\\ &<\epsilon \end{align*}\] (1) rewrite the expression (2) triangle inequality (3) we know the individual limits exist if we choose to take the least of the two \(\delta\)s for the known limits. In other words, we're given \(0<|x-c|<\delta_1\) for \(f\) and \(0<|x-c|<\delta_2\) for \(g\) - we pick \(\delta=\min\{\delta_1,\delta_2\}\).

OpenStudy (anonymous):

Thank you!

OpenStudy (xapproachesinfinity):

@SithsAndGiggles, thanks for the proof can we say a ==> b and c==>d then a+c ==>b+d

OpenStudy (xapproachesinfinity):

@ganeshie8

OpenStudy (tkhunny):

Have to object to Siths very clean demonstration. The very first statement is that the limits exist. Without the assumption that the limits exist, it is an entirely different story. Please consider the two functions I suggested and don't make the assumption of existing limits.

OpenStudy (xapproachesinfinity):

Well he assumed the limits exist in the first line too? didn't he or may be im understanding it in a wrong way.Thanks for checking this

OpenStudy (tkhunny):

That is what I said, "The very first statement is that the limits exist." Well, it's not perfectly obvious that this is the first statement, but it is clear later that it is the intent. A perfectly clear statement would have been: \(\lim\limits_{x\rightarrow c}\;f(x) = L < \infty\) This says rather deliberately that the limit exists and is finite.

OpenStudy (anonymous):

@sithsandgiggles Are you sure that this would work for proving the converse of lim (f(x) + g(x)) = f(x) + g(x). It looks like the straightforward proof that I know. I know what a converse is, I'm just clueless on setting up a proof for it. I also have to prove the converse of lim(f(x)*g(x)) = limf(x)* limg(x). Thanks for your help!

OpenStudy (larseighner):

The statement does not have a converse because it is not an implication. It is an equality. And an equality works both ways.

OpenStudy (tkhunny):

It also doesn't have a converse because it isn't true at all as it is stated. Counter examples are trivially observed.

OpenStudy (anonymous):

Everything you need is available here: http://tutorial.math.lamar.edu/Classes/CalcI/LimitProofs.aspx Notice the language used in the statements proved in the link.

OpenStudy (anonymous):

Thanks so much for your help guys! I get what you all are saying, I'm just confused more about the question itself.

OpenStudy (tkhunny):

Really, it is this: 1) If you have a counter example, you are done. 2) If you construct a valid proof, you are done. 3) Additional assumptions may be added in order to gain greater insight and to prove lesser statements. 4) Fewer assumptions are the sort of thing that advanced studies are about - The Generalization! It is generally agreed that a generalization should have at least two examples. :-)